OpenStudy (anonymous):

what is x in log(6x+5)-log2=log5

OpenStudy (alekos):

any ideas? have you tried this?

OpenStudy (anonymous):

separate the logs first

OpenStudy (anonymous):

log 6x+log 5-log2=log5

OpenStudy (anonymous):

bring over log 5 to right side cancels

OpenStudy (anonymous):

log 6x=log2

OpenStudy (anonymous):

log 6 +log x= log 2

OpenStudy (anonymous):

logx=log2-log6

OpenStudy (anonymous):

logx=log(2/6)

OpenStudy (anonymous):

logx=log(1/3)

OpenStudy (alekos):

have you finished?

OpenStudy (anonymous):

no haven't finished yet

OpenStudy (alekos):

so whats x?

OpenStudy (anonymous):

x=10^(log(1/3))

OpenStudy (alekos):

which is?

OpenStudy (alekos):

If you're method was correct then x=1/3 what does your answer book say?

OpenStudy (anonymous):

i get 1/3 when i punch it into the calculator

OpenStudy (anonymous):

what answer book :S

OpenStudy (alekos):

dont you have answers at the back of the book?

OpenStudy (anonymous):

I'm just doing these questions off the top of my head i'm a university student :P

OpenStudy (alekos):

wow, are you doing uni maths?

OpenStudy (anonymous):

yup i miss high-school and elementary school math its easy compared to what i have to do now T.T

OpenStudy (alekos):

nice try, but the method is incorrect. log laws 1. log(ab) = loga + logb 2. log(a/b) = loga - logb 3. if loga = logb, then a=b

OpenStudy (alekos):

so for log(6x+5) - log2 = log5 we have log[(6x+5)/2] = log5 therefore (6x+5)/2 = 5 and 6x+5 = 10 6x = 5 hence x = 5/6

OpenStudy (anonymous):

i need to brush up on these old skills

OpenStudy (alekos):

yes, you do

OpenStudy (alekos):

if you're doing uni maths you need to know your basics inside out

OpenStudy (anonymous):

trust me u lose a couple of things down the road cuz i hardly touched upon logs and when u look back its way diff

OpenStudy (alekos):

i understand. same thing happened to me when i did uni maths. i think it was series at the time

OpenStudy (anonymous):

I'm in good hands then this website is going to be perfect to answering any math questions i may have :D

OpenStudy (alekos):

yeah its a pretty good site

OpenStudy (alekos):

thanks for the medal